# PIM-malloc: A Fast and Scalable Dynamic Memory Allocator for Processing-In-Memory (PIM) Architectures

### v2

@@ 3.2. Design Considerations for Dynamic Memory Allocators in PIM
In Figure 5(a), we show how memory allocation latency changes as the number of PIM cores requesting allocations increases. Notably, with the exception of “PIM-Metadata/PIM-Executed” (red), allocation latency increases for all other design points as the number of PIM cores grows. Both “Host-Metadata/PIM-Executed” (black) and “PIM-Metadata/Host-Executed” (yellow) designs require metadata to be transferred to the processors executing the buddy allocation algorithm, involving explicit data transfers between standard DRAM and PIM memory. Since metadata size scales with the number of PIM cores, the data transfer overhead becomes substantial, resulting in increased system-wide memory allocation latency (Figure 5(b)). Although “Host-Metadata/Host-Executed” (gray) avoids excessive data movement between DRAM and PIM, it demonstrates limited scalability due to CPU’s limited parallel processing capability. For example, with 512 PIM cores in our baseline UPMEM-PIM, there may be up to 512 independent memory allocation requests that can be serviced in parallel. However, the limited number of CPU cores constrains the number of memory allocation requests that can be handled concurrently, reducing system-wide performance. In contrast, despite relying on wimpy PIM cores, “PIM-Metadata/PIM-Executed” (red) provides a highly scalable solution. This is because “PIM-Metadata/PIM-Executed” eliminates data transfer overhead between DRAM and PIM while allowing each PIM core to handle its allocation requests locally and in parallel. Overall, we conclude that the “PIM-Metadata/PIM-Executed” approach is the optimal baseline design point for implementing a high-performance PIM memory allocator. For the remainder of this paper, both the straw-man buddy_alloc_PIM_DRAM and our proposed PIM-malloc design are based on this design principle.
3.3. Remaining Challenges with the Straw-man PIM Buddy Allocator
@@
While our design space exploration in Section 3.2 identified the “PIM-Metadata/PIM-Executed” approach as a strong foundation for the straw-man buddy_alloc_PIM_DRAM design222 For brevity, this paper henceforth interchangeably refers to the “PIM-Metadata/PIM-Executed” straw-man buddy_alloc_PIM_DRAM as the straw-man PIM buddy allocator., challenges still remain that must be addressed to develop a fast and scalable PIM memory allocator. We detail the two limitations of the straw-man PIM buddy allocator.
@@
First, the inherent design of the buddy allocator causes latency to increase when the heap size expands or the minimum (de)allocation size decreases. The buddy allocator manages the heap using a tree data structure, where larger heap sizes or smaller (de)allocation sizes increase the tree’s depth, requiring deeper tree traversals. This results in longer metadata search times and higher latency during memory (de)allocations. In UPMEM-PIM’s existing scratchpad-based buddy allocator (buddy_alloc()), the maximum heap size in the scratchpad (WRAM) is capped at 64 KB, with the minimum (de)allocation size set to 32 bytes, resulting in a maximum of a 10-level buddy tree traversal (). Consequently, the baseline scratchpad-based buddy allocator’s memory allocation latency is sufficiently low enough to avoid becoming a performance bottleneck. With our straw-man PIM buddy allocator, however, the heap size can be orders of magnitude larger than the PIM core’s scratchpad space, leading to a maximum 20-level tree traversal with heap size of 32 MB and causing significant increases in memory allocation latency. Figure 6 compares the average memory allocation latency of our straw-man PIM buddy allocator over various heap sizes and minimum memory (de)allocation sizes. Compared to a 2 KB allocation in a 32 KB heap, a 32 B allocation over a 32 MB heap experiences up to a 12 slowdown, highlighting the scalability challenges of our straw-man PIM buddy allocator.
@@
The second challenge of the straw-man PIM buddy allocator is its susceptibility to thread contention. The UPMEM-PIM programming model employs fine-grained multithreading within each PIM core (DPU) to achieve high performance. Ideally, the PIM memory allocator should provide low-latency, high-throughput allocation services even under such multithreaded environment. Unfortunately, when multiple threads simultaneously request (de)allocation within a given timeframe, the buddy allocator – protected by a mutex – can only process one thread’s allocation and deallocation request at a time, forcing the remaining threads to be stuck in a busy-waiting state. In Figure 7, we show the effect of thread contention on memory allocation performance. Under a 16-thread execution scenario, severe thread contention causes significant fluctuations in memory allocation latency, unlike in single-thread memory allocation where the latency remains fairly stable as shown in Figure 7(a). This behavior occurs because it is busy-waiting for its turn to acquire the mutex and have its allocation request serviced (Figure 7(b)).
4. PIM-malloc: A Fast and Scalable Dynamic Memory Allocator for PIM
@@
This section details the design of PIM-malloc, our fast and scalable dynamic memory allocator for modern bank-level PIM architectures. We present two different versions of PIM-malloc: a software-only implementation, henceforth referred to as PIM-malloc-SW, and a hardware/software co-design, henceforth referred to as PIM-malloc-HW/SW. PIM-malloc-SW effectively addresses the limitations of the straw-man PIM buddy allocator purely through software optimizations, allowing it to be readily deployable on existing PIM device. To achieve further performance gains, we propose PIM-malloc-HW/SW, a hardware/software co-design incorporating lightweight modifications to the PIM core.
4.1. Software-only PIM-malloc-SW
@@
Design principle. The design of PIM-malloc-SW is motivated by the need to address limitations in the straw-man PIM buddy allocator. There are two key challenges associated with our straw-man allocator design: (1) buddy allocators are prone to thread contention, and (2) managing the heap area using only the buddy allocation algorithm results in increased latency, especially when the heap size is large or allocation/deallocation request sizes are small. To address these issues, PIM-malloc-SW employs a hierarchical memory allocation scheme, as illustrated in Figure 8(a): (1) At the frontend, a per-thread thread cache manages small memory allocation and deallocation requests (between 16 bytes and 2 KB). (2) For allocation requests that exceed this limit, PIM-malloc-SW employs the buddy allocator at the backend to handle allocation sizes exceeding 2 KB.
@@
In PIM-malloc-SW, thread caches are memory pools exclusively assigned to each thread. Since each PIM core in UPMEM-PIM can support up to 24 concurrent threads, PIM-malloc-SW allows for up to 24 thread caches per PIM core (Figure 8(b)). This design enables each thread to efficiently manage small memory (de)allocations independently, without requiring mutex acquisition, thereby eliminating inter-thread contention. While PIM-malloc-SW’s buddy allocator still uses a shared mutex for multi-threaded allocations and does not support exclusive per-thread memory pools, it benefits from reduced lock contention due to our hierarchical memory allocator design. Specifically, the frontend memory allocator (thread cache) can filter out allocation requests, reducing the load on the backend memory allocator (buddy allocator). Naturally, the effectiveness of the thread cache’s request filtering depends on the target application’s memory allocation patterns. Prior work has observed that real-world application’s memory allocation sizes are generally concentrated in small requests. For example, allocation requests under 1 KB account for 98% of datacenter allocations (Kanev et al., 2017; Hunter et al., 2021; Zhou et al., 2024), and those under 512 B comprise 93% of serverless allocations (Wang et al., 2023b), justifying our hierarchical allocator design. By handling these small sized requests locally within thread caches, contention on the shared buddy allocator is minimized, reducing congestion and lowering busy-wait latency. Another key advantage of our hierarchical design is its ability to reduce the buddy allocator’s metadata overhead compared to the single-level buddy allocation algorithm used in the straw-man PIM buddy allocator, which handles both small and large allocation requests. Reducing the metadata tree depth from 20 () to 13 () substantially decreases the buddy allocator’s tree traversal overhead. This improvement enhances (de)allocation speeds for large memory allocations.
@@
Since PIM-malloc-SW employs a hierarchical allocator design with private thread caches and a shared buddy allocator, one might wonder why we limit our allocator hierarchy to only two layers. Why not adopt a deeper hierarchical allocator design, like TCMalloc (Ghemawat and Menage, 2009), which features four allocator layers and incorporates more sophisticated heap management techniques, such as work-stealing from other private CPU caches and prefetching memory blocks across different layers? Although this approach might seem advantageous from a performance standpoint, adapting the complex memory allocators used in TCMalloc to the PIM environment is impractical for several reasons, as detailed below.
First, current PIM devices are fabricated using density-optimized technology nodes (e.g., 20 nm DRAM technology for UPMEM-PIM (Devaux, 2019)), which impose significant design constraints. These constraints prevent the use of advanced microarchitectures and large on-chip SRAM. For example, PIM devices have highly limited on-chip storage for program binaries, with UPMEM-PIM offering a maximum instruction memory (IRAM) size of only 24 KB to store PIM program instructions. Implementing a complex memory allocator like TCMalloc, which consists of approximately 60,000 lines of C++ code, is therefore infeasible within these constraints.
@@
Additionally, current PIM systems lack essential system software support like dynamic thread launch, further hindering the implementation of sophisticated memory allocators like TCMalloc. Given these challenges, our design aims for a streamlined and efficient approach tailored to the unique resource constraints of PIM devices.
@@
Implementation. Figure 8(b) details the implementation of the thread cache and its interaction with the buddy allocator. The thread cache manages small-sized memory (de)allocations using 4 KB memory blocks provided by the buddy allocator. When free memory blocks are exhausted and additional memory is needed, the thread cache issues allocation requests to the buddy allocator. Conversely, when all memory sub-blocks within a 4 KB block in the thread cache are freed by the PIM program, they are merged back into a 4 KB block and returned to the buddy allocator.
In our design, each thread cache contains eight linked lists to efficiently manage memory (de)allocations of specific sizes. For instance, the linked list at index 0 handles 16 B (de)allocations, while the linked list at index 7 manages 2 KB (de)allocations. Each linked list can hold multiple 4 KB memory blocks received from the buddy allocator, which are divided into smaller sub-blocks managed by that list (e.g., a single 4 KB block is divided into two 2 KB sub-blocks in the linked list at index 7). This sub-block size, termed the size class, ensures efficient management and allocation of small memory requests. By maintaining a pool of fixed-size sub-blocks at the private thread caches, PIM-malloc-SW can quickly fulfill small-sized memory allocation requests with O(1) latency. To track the allocation status of each sub-block in thread caches, we assign a dedicated 1-bit metadata per sub-block. The buddy allocator is implemented identically to the straw-man PIM buddy allocator described in Section 3.2, featuring a software-managed metadata buffer to cache recently accessed buddy allocator metadata – except that the buddy allocator’s tree depth is reduced from 20 to 13.
Software interface. PIM-malloc-SW provides UPMEM-PIM compatible APIs that enable PIM programmers to initialize and manage heap memory (Table 2). Before running a PIM program, PIM-malloc-SW requires initialization, which involves several key steps. First, it resets metadata within both the thread caches and the buddy allocator to ensure a clean and consistent state for memory management. Additionally, the initialization process pre-populates the thread caches with free memory blocks (a single 4 KB block for each linked list within the thread cache), allowing immediate memory allocation when requested. This proactive strategy reduces initial allocation latency and improves system responsiveness. Since initialization is a one-time operation, a designated thread (i.e., the thread with the thread ID of ‘0’) handles this process. After initialization, each thread can independently use pimMalloc() and pimFree().
@@
(Case #3: thread cache bypass) This scenario involves an allocation request exceeding the thread cache’s largest size class, i.e., the sub-block size. For example, when a 8 KB block is requested, the allocator bypasses the thread cache entirely, as its operational range is limited to allocations between 16 bytes and 2 KB. Instead, PIM-malloc-SW directly queries the buddy allocator for the 8 KB allocation ( in Figure 9(c)). The result from the buddy allocator is returned back to the user program to fulfill the memory allocation request ( ).
4.2. Hardware/Software Co-Designed PIM-malloc-HW/SW
@@
Limitations of PIM-malloc-SW. While PIM-malloc-SW was able to achieve a remarkable speedup over the straw-man PIM buddy allocator, the unique hardware/software constraints of the PIM system pose significant challenges for further improving PIM-malloc-SW via software-only optimizations. As an example, consider PIM-malloc-SW’s software-managed metadata buffer which is used to cache recently accessed metadata and exploit locality. For the simplicity of its design, the baseline implementation of PIM-malloc-SW employs a coarse-grained metadata buffer management strategy where a miss at the software-managed buffer leads to flushing this buffer and a subsequent reloading of the required metadata from DRAM. To address this inefficiency, we also explored a fine-grained metadata buffer management mechanism based on an LRU replacement policy, implemented entirely in software. However, this approach encountered several challenges stemming from the inherent computational limitations of PIM. Specifically, such fine-grained buffer management introduces significant software overhead, negating the benefits of reduced DRAM transfers via intelligent caching decisions. As such, our analysis reveals that, while the aforementioned LRU-based metadata buffer management effectively reduces metadata transfers from DRAM, the associated computational overhead results in negligible or even diminished performance gains, which led to our baseline, coarse-grained metadata buffer management in PIM-malloc-SW. Overall, the difficulty of further improving PIM-malloc-SW’s performance purely via software-level optimizations motivates our hardware/software co-design approach, aka PIM-malloc-HW/SW, detailed below.
@@
Prior work on HW-accelerated memory allocators and its differences vs. PIM-malloc-HW/SW. Several studies have explored hardware-based approaches to optimize dynamic memory allocators in CPUs (Maas et al., 2018; Chang and Gehringer, 1996; Cam et al., 1999; Li et al., 2006; Von Puttkamer, 1975; Kanev et al., 2017; Wang et al., 2023b). Notably, Mallacc (Kanev et al., 2017) augments the processor core with a dedicated hardware cache to enhance the performance of TCMalloc’s hierarchical memory allocator design. This work primarily targets the frontend memory allocator (per-CPU cache) of TCMalloc, which is based on the observation that most allocation requests can be satisfied at this layer. One might wonder whether directly applying the approaches suggested in Mallacc (Kanev et al., 2017) to PIM-malloc, which is to accelerate PIM-malloc-SW’s frontend memory allocator, could yield performance benefits. However, our characterization of PIM-malloc-SW revealed unique properties in its two-layer PIM memory allocator design that necessitate a fundamentally different approach than what was taken in Mallacc (Kanev et al., 2017).
In Figure 10, we analyze the memory allocation patterns of a dynamic graph update application (Section 5 provide details of this workload), which provides several key insights. First, more than 90% of memory allocation requests are serviced by PIM-malloc-SW’s frontend (Figure 10(a)). Second, allocation requests that cannot be serviced at the frontend, which fall back to the backend, incur approximately 80 higher per-allocation latency compared to the frontend (Figure 10(b)). TCMalloc employs a four-layer hierarchical memory allocator structure with advanced memory pool management techniques, such as memory block prefetching. Consequently, the performance difference between TCMalloc’s frontend and backend (aka TransferCache) allocators is “only” 20 (Zhou et al., 2024), which is moderate than PIM-malloc-SW’s 80 performance gap. Therefore, third, when evaluated with our dynamic graph update workload, we observe that 87% of the aggregate PIM memory allocation latency originates from the backend allocator (Figure 10(c)), which is primarily due to its significantly higher per-allocation latency compared to the frontend allocator. This is in stark contrast to TCMalloc, which consumes 53% of its memory allocation time in the frontend memory allocator (Zhou et al., 2024), unlike PIM-malloc-SW’s frontend allocator which consumes only 13% (Figure 10(c)).
Thus, we design PIM-malloc-HW/SW to improve the performance of the backend, unlike prior approach (Kanev et al., 2017) which focuses on accelerating the top-most memory allocator.
@@
The buddy cache is implemented as a fully-associative cache using Content-Addressable Memory (CAM). Each entry consists of a valid bit (1 bit), the DRAM address of the metadata (4 bytes) serving as the tag, and the metadata value (4 bytes). Our baseline buddy cache utilizes a 16-entry CAM structure (4 B metadata per entry, 16 entries in total) because storing 64 B of metadata per cache was sufficient to capture most access locality (Figure 15 details sensitivity to this design parameter). The buddy cache is managed using an LRU replacement policy for efficient metadata retrieval and eviction. PIM-malloc-HW/SW extends the Instruction Set Architecture (ISA) of our PIM design with the following four instructions to interface the buddy cache with the PIM-malloc runtime system:
@@
HW/SW interface for the buddy cache. Figure 12 shows code snippets of the getMetadata() function used as part of implementing the buddy allocator in both PIM-malloc-SW (Figure 12(a)) and PIM-malloc-HW/SW (Figure 12(b)). This function retrieves the metadata of a given tree node (metadataIdx) accessed during buddy tree traversal. Initially, both implementations search for the requested metadata in their respective on-chip SRAM storage: (1) PIM-malloc-SW checks the software-managed metadata buffer stored inside the scratchpad (searchBuddyBuffer()), while (2) PIM-malloc-HW/SW checks the hardware-based buddy cache (lookup_buddy_cache()) by using the lookup_bc instruction. In the case of a cache hit (case 1, line 2), both implementations return the metadata directly from their respective caches, with PIM-malloc-HW/SW utilizing the read_bc instruction as part of the execution of the read_buddy_cache() function. However, cache misses (case 2, line 6) can also occur which trigger distinct handling procedures. In PIM-malloc-SW, the software-managed metadata buffer is flushed (Figure 12(a), line 9) and the required metadata is fetched from DRAM (Figure 12(a), line 12) to populate the metadata buffer with both the requested and adjacent metadata (Figure 12(a), line 15). As for PIM-malloc-HW/SW, the hardware-based buddy cache facilitates the implementation of a more fine-grained and sophisticated cache replacement policy. Specifically, upon a cache miss, PIM-malloc-HW/SW fetches only the requested metadata from DRAM (Figure 12(b), line 9) and updates the buddy cache by evicting just the LRU entry using the write_bc instruction (Figure 12(b), line 13). Compared to PIM-malloc-SW, this approach helps not only minimize data transfers associated with metadata retrieval, but it also increases the buddy cache hit rates.

### v3

@@ III-B Design Space of Dynamic Memory Allocators in PIM
In Figure 6(a), we show how memory allocation latency changes as the number of PIM cores requesting allocations increases. Notably, with the exception of “PIM-Metadata/PIM-Executed” (red), allocation latency increases for all other design points as the number of PIM cores grows. Both “Host-Metadata/PIM-Executed” (black) and “PIM-Metadata/Host-Executed” (yellow) designs require metadata to be transferred to the processors executing the buddy allocation algorithm, involving explicit data transfers between standard DRAM and PIM memory. Since metadata size scales with the number of PIM cores, the data transfer overhead becomes substantial, resulting in increased system-wide memory allocation latency (Figure 6(b)). Although “Host-Metadata/Host-Executed” (gray) avoids excessive data movement between DRAM and PIM, it demonstrates limited scalability due to the CPU’s constrained parallelism. In contrast, despite relying on wimpy PIM cores, “PIM-Metadata/PIM-Executed” (red) provides a highly scalable solution. This is because “PIM-Metadata/PIM-Executed” eliminates data transfer overhead between DRAM and PIM while allowing each PIM core to handle its allocation requests locally and in parallel. Overall, we conclude that the “PIM-Metadata/PIM-Executed” approach is the optimal baseline design point for implementing a high-performance PIM memory allocator. For the remainder of this paper, both the straw-man buddy_alloc_PIM_DRAM and our proposed PIM-malloc design are based on this design principle.
III-C Remaining Challenges with the Straw-man PIM Allocator
@@
While our design space exploration in Section III-B identified the “PIM-Metadata/PIM-Executed” approach as a strong foundation for the straw-man buddy_alloc_PIM_DRAM design222We interchangeably refer to the “PIM-Metadata/PIM-Executed” straw-man buddy_alloc_PIM_DRAM as straw-man PIM buddy allocator for brevity., challenges still remain that must be addressed to develop a fast and scalable PIM memory allocator. We detail the two limitations of the straw-man PIM buddy allocator.
@@
First, the inherent design of the buddy allocator causes latency to increase when the heap size expands or the minimum (de)allocation size decreases. The buddy allocator manages the heap using a tree data structure, where larger heap sizes or smaller (de)allocation sizes increase the tree’s depth, requiring deeper tree traversals. This results in longer metadata search times and higher latency during memory (de)allocations. In UPMEM-PIM’s existing scratchpad-based buddy allocator (buddy_alloc()), the maximum heap size in the scratchpad (WRAM) is capped at 64 KB, with the minimum (de)allocation size set to 32 bytes, resulting in a maximum of a 10-level buddy tree traversal (). Consequently, the baseline scratchpad-based buddy allocator’s memory allocation latency is low enough to avoid becoming a performance bottleneck. With our straw-man PIM buddy allocator, however, the heap size can be orders of magnitude larger than the PIM core’s scratchpad space, leading to a maximum 20-level tree traversal with heap size of 32 MB and causing significant increases in memory allocation latency. Figure 7 compares the average memory allocation latency of our straw-man PIM buddy allocator over various heap sizes and minimum memory (de)allocation sizes. Compared to a 2 KB allocation in a 32 KB heap, a 32 B allocation over a 32 MB heap experiences up to a 12 slowdown, highlighting the scalability challenges of our straw-man PIM buddy allocator.
@@
The second challenge of the straw-man PIM buddy allocator is its susceptibility to thread contention. The UPMEM-PIM programming model employs fine-grained multithreading within each PIM core (DPU) to achieve high performance. When multiple threads simultaneously request (de)allocation within a given timeframe, the buddy allocator – protected by a mutex – can only process one thread’s allocation and deallocation request at a time, forcing the remaining threads to be stuck in a busy-waiting state. In Figure 8, we show the effect of thread contention on memory allocation performance. Under a 16-thread execution scenario, severe thread contention causes significant fluctuations in memory allocation latency, unlike in single-thread memory allocation where the latency remains fairly stable as shown in Figure 8(a). This behavior occurs because it is busy-waiting for its turn to acquire the mutex and have its allocation request serviced (Figure 8(b)).
IV PIM-malloc: A Fast and Scalable Dynamic Memory Allocator for PIM
@@
We present two different versions of PIM-malloc. PIM-malloc-SW effectively addresses the limitations of the straw-man PIM buddy allocator purely through software optimizations, allowing it to be readily deployable on existing PIM devices. To achieve further performance gains, we propose PIM-malloc-HW/SW, a hardware/software co-design incorporating lightweight modifications to the PIM core.
IV-A Software-only PIM-malloc-SW
@@
Design principle. To address limitations in the straw-man PIM buddy allocator, PIM-malloc-SW employs a hierarchical memory allocation scheme, as illustrated in Figure 9(a): (1) At the frontend, a per-thread thread cache manages small memory (de)allocation requests (between 16 bytes and 2 KB). (2) For allocation requests that exceed 2 KB, PIM-malloc-SW employs the buddy allocator in the backend.
@@
In PIM-malloc-SW, thread caches are memory pools exclusively assigned to each thread. Since each PIM core in UPMEM-PIM can support up to 24 concurrent threads, PIM-malloc-SW allows for up to 24 thread caches per PIM core (Figure 9(b)). This design enables each thread to efficiently manage small memory (de)allocations independently, without requiring mutex acquisition. While PIM-malloc-SW’s buddy allocator still uses a shared mutex for multi-threaded allocations and does not support exclusive per-thread memory pools, it benefits from reduced lock contention due to our hierarchical memory allocator design. Specifically, the frontend allocator can filter out allocation requests, reducing the load on the backend allocator. Another key advantage of our hierarchical design is its ability to reduce the buddy allocator’s metadata overhead compared to the single-level buddy allocation algorithm used in the straw-man PIM buddy allocator, which handles both small and large allocation requests. Reducing the metadata tree depth from 20 () to 13 () substantially decreases the buddy allocator’s tree traversal overhead. This improvement enhances (de)allocation speeds for large memory allocations.
@@
Since PIM-malloc-SW employs a hierarchical allocator design with private thread caches and a shared buddy allocator, one might wonder why we limit our allocator hierarchy to only two layers. Why not adopt a deeper hierarchical allocator design, like TCMalloc [40], which features four allocator layers and incorporates more sophisticated heap management techniques, such as prefetching memory blocks across different layers? Although this approach might seem advantageous from a performance standpoint, adapting the complex memory allocators to the PIM is impractical for several reasons.
@@
First, PIM devices are fabricated on DRAM process (e.g., 20 nm DRAM for UPMEM-PIM [29]), which imposes severe constraints on their microarchitecture and on-chip storage. For instance, the mere 24 KB of instruction memory (IRAM) in UPMEM-PIM renders complex memory allocators like TCMalloc (approximately 60K C++ lines) infeasible. In contrast, PIM-malloc, with only about 1,000 lines of code, is well-suited for the tight resource budget of PIM devices. Additionally, current PIM systems lack essential system software support like dynamic thread launch, further hindering the implementation of sophisticated memory allocators. Given these challenges, our design aims for a streamlined and efficient approach tailored to the unique resource constraints of PIM devices.
Implementation. Figure 9(b) details the implementation of the thread cache and its interaction with the buddy allocator. The thread cache manages small-sized memory (de)allocations using 4 KB memory blocks provided by the buddy allocator. When free memory blocks are exhausted, the thread cache issues allocation requests to the buddy allocator. Conversely, when all memory sub-blocks within a 4 KB block in the thread cache are freed by the PIM program, they are merged back into a 4 KB block and returned to the buddy allocator.
In our design, each thread cache contains eight linked lists to efficiently manage memory (de)allocations of specific sizes. For instance, the linked list at index 0 handles 16 B (de)allocations, while the linked list at index 7 manages 2 KB (de)allocations. Each linked list can hold multiple 4 KB memory blocks received from the buddy allocator, which are divided into smaller sub-blocks managed by that list (e.g., a single 4 KB block is divided into two 2 KB sub-blocks in the linked list at index 7). This sub-block size, termed the size class, ensures efficient management and allocation of small memory requests. By maintaining a pool of fixed-size sub-blocks at the private thread caches, PIM-malloc-SW can quickly fulfill small-sized memory allocation requests with O(1) latency. A key benefit of this size class-based approach is the elimination of external fragmentation within the thread caches. By design, each linked list functions as an independent pool of fixed-size memory chunks. The state of one pool (for instance, a large number of free 32 B chunks) therefore has no bearing on requests for another, such as a 1 KB allocation, because they are serviced by entirely separate pools. This design choice effectively mitigates the fragmentation problem within the thread caches. To track the allocation status of each sub-block in thread caches, we assign dedicated-bit metadata for each sub-block. As shown in Figure 9(b), each of the eight linked lists maintains its own metadata bitmap (yellow-colored box) to track the allocation status of its sub-blocks. For example, to handle a 16 B allocation request, PIM-malloc-SW searches the metadata bitmap at index 0 for a bit indicating an available sub-block (i.e., a bit with a value of 1). Once it finds such a bit, PIM-malloc-SW calculates and returns the address of the corresponding 16 B sub-block (red-colored box) based on the bit’s position (offset). Similarly, a 2 KB request is processed using the metadata bitmap at index 7. The buddy allocator is implemented identically to the straw-man PIM buddy allocator described in Section III-B, featuring a software-managed metadata buffer to cache recently accessed buddy allocator metadata – except that the buddy allocator’s tree depth is reduced from 20 to 13.
@@
Software interface. PIM-malloc-SW provides UPMEM-PIM compatible APIs that enable PIM programmers to initialize and manage heap memory (Table II). Before running a PIM program, PIM-malloc-SW requires initialization, which involves several key steps. First, it resets the metadata of the allocator to ensure a clean and consistent state for memory management. Additionally, the initialization process pre-populates the thread caches with free memory blocks (a single 4 KB block for each linked list within the thread cache), allowing immediate memory allocation when requested. This proactive strategy reduces initial allocation latency and improves system responsiveness. Since initialization is a one-time operation, a designated thread (i.e., the thread with the thread ID of ‘0’) handles this process. After initialization, each thread can independently use pimMalloc() and pimFree().
@@
(Case #3: thread cache bypass) This scenario involves an allocation request exceeding the thread cache’s largest size class, i.e., the sub-block size. For example, when a 8 KB block is requested, the allocator bypasses the thread cache entirely, as its operational range is limited to allocations between 16 bytes and 2 KB. Instead, PIM-malloc-SW directly queries the buddy allocator for the 8 KB allocation ( in Figure 10(c)). The result from the buddy allocator is returned to the user program to fulfill the memory allocation request ( ).
IV-B Hardware/Software Co-Designed PIM-malloc-HW/SW
@@
Limitations of PIM-malloc-SW. While PIM-malloc-SW was able to achieve a remarkable speedup over the straw-man PIM buddy allocator, the unique hardware/software constraints of the PIM system pose significant challenges for further improving PIM-malloc-SW via software-only optimizations. As an example, consider PIM-malloc-SW’s software-managed metadata buffer which is used to cache recently accessed metadata and exploit locality. For the simplicity of its design, the baseline implementation of PIM-malloc-SW employs a coarse-grained metadata buffer management strategy where a miss at the software-managed buffer leads to flushing this buffer and a subsequent reloading of the required metadata from DRAM. To address this inefficiency, we also explored a fine-grained metadata buffer management mechanism based on an LRU replacement policy, implemented entirely in software. However, this approach encountered several challenges stemming from the inherent computational limitations of PIM. Specifically, such fine-grained buffer management introduces significant software overhead, negating the benefits of reduced DRAM transfers via intelligent caching decisions. As such, our analysis reveals that, while the aforementioned LRU-based metadata buffer management effectively reduces metadata transfers from DRAM, the associated computational overhead results in negligible or even diminished performance gains (fine‑grained buffer management shows a 29% performance degradation in the microbenchmark described in Section V, where 16 threads request 4 KB allocations), which led to our baseline, coarse-grained metadata buffer management in PIM-malloc-SW. Overall, the difficulty of further improving PIM-malloc-SW’s performance purely via software-level optimizations motivates our hardware/software co-design approach, aka PIM-malloc-HW/SW, detailed below.
@@
Prior work on HW-accelerated memory allocators and its differences vs. PIM-malloc-HW/SW. Several studies have explored hardware-based approaches to optimize dynamic memory allocators in CPUs [94, 18, 17, 91, 131, 68, 133]. Notably, Mallacc [68] augments the processor core with a dedicated hardware cache to enhance the performance of TCMalloc’s hierarchical memory allocator design. This work primarily targets the frontend memory allocator (per-CPU cache) of TCMalloc, which is based on the observation that most allocation requests can be satisfied at this layer. One might wonder whether directly applying the approaches suggested in Mallacc [68] to PIM-malloc, which aims to accelerate PIM-malloc-SW’s frontend memory allocator, could yield performance benefits. However, our characterization of PIM-malloc-SW revealed unique properties in its two-layer PIM memory allocator design that necessitate a fundamentally different approach than what was taken in Mallacc [68].
In Figure 11, we analyze memory allocation behavior for our workloads (see Section V for its implementation details) and provide several key insights. First, on average, 93% of allocation requests are satisfied by PIM-malloc-SW’s frontend (Figure 11(a)). Second, although the backend allocator serves far fewer requests, it accounts for 68% of the total allocation latency (Figure 11(b)). This is in stark contrast to TCMalloc, which consumes 53% of its memory allocation time in the frontend memory allocator [142], unlike PIM-malloc-SW’s frontend allocator which contributes only 32% (Figure 11(b)). Such phenomenon arises from differences between PIM-malloc and TCMalloc. TCMalloc employs a four‑layer hierarchical memory allocator with advanced memory‑pool management techniques, such as memory‑block prefetching, whereas PIM-malloc lacks these features because of PIM constraints, as discussed in Section IV-A. Thus, we design PIM-malloc-HW/SW to improve the performance of the backend allocator, unlike the prior approach [68] which focuses on accelerating the top-level memory allocator.
@@
The buddy cache is implemented as a fully-associative cache using Content-Addressable Memory (CAM). Each entry consists of a valid bit (1 bit), the DRAM address of the metadata (4 bytes) serving as the tag, and the metadata value (4 bytes). Our baseline buddy cache utilizes a 16-entry CAM structure (4 B metadata per entry, 16 entries in total) because storing 64 B of metadata per cache was sufficient to capture most access locality (Figure 16 details sensitivity to this design parameter). The buddy cache is managed using an LRU replacement policy for efficient metadata retrieval and eviction. PIM-malloc-HW/SW extends the Instruction Set Architecture (ISA) of our PIM design with the following four instructions to interface the buddy cache with the runtime system:
@@
HW/SW interface for the buddy cache. Figure 13 shows code snippets of the getMetadata() function used as part of implementing the buddy allocator in both PIM-malloc-SW (Figure 13(a)) and PIM-malloc-HW/SW (Figure 13(b)). This function retrieves the metadata of a given tree node (metadataIdx) accessed during buddy tree traversal. Initially, both implementations search for the requested metadata in their respective on-chip SRAM storage: (1) PIM-malloc-SW checks the software-managed metadata buffer stored inside the scratchpad (searchBuddyBuffer()), while (2) PIM-malloc-HW/SW checks the hardware-based buddy cache (lookup_buddy_cache()) by using the lookup_bc instruction. In the case of a cache hit (case 1, line 2), both implementations return the metadata directly from their respective caches, with PIM-malloc-HW/SW utilizing the read_bc instruction as part of the execution of the read_buddy_cache() function. However, cache misses (case 2, line 6) can also occur which trigger distinct handling procedures. In PIM-malloc-SW, the software-managed metadata buffer is flushed (Figure 13(a), line 9) and the required metadata is fetched from DRAM (Figure 13(a), line 12) to populate the metadata buffer with both the requested and adjacent metadata (Figure 13(a), line 15). As for PIM-malloc-HW/SW, the hardware-based buddy cache facilitates the implementation of a more fine-grained and sophisticated cache replacement policy. Specifically, upon a cache miss, PIM-malloc-HW/SW fetches only the requested metadata from DRAM (Figure 13(b), line 9) and updates the buddy cache by evicting just the LRU entry using the write_bc instruction (Figure 13(b), line 13). Compared with PIM-malloc-SW, this approach not only minimizes metadata related data transfers but also increases the buddy cache hit rate.